The average molar mass of a mixture of methane and ethene present in the ratio a:b is found to be 20 g/mol.  If the ration were reversed, what would be the molar mass of the mixture? 

Grade:11

2 Answers

Arun Kumar IIT Delhi
9 years ago

Arun Kumar
IIT Delhi
Askiitians Faculty
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Molar mass of methane = 16 gmol^-1
Molar mass of ethane = 28 gmol^-1
When they are present in the a:b their average molar mass
= (ax16 + bx28)/(a+b) = 20
i.e., 16a + 28b = 20(a+b)
or 4a + 7b = 5 (a+b)
or a = 2b
or a/b = 2/1 = 2:1
If the ratio is reversed, now the ratio
a:b = 1:2
Average molar mass = 1x16 + 2x28/1+2
= 24 gmol^-1

Thanks and Regards